If the perimeter is 32 feet and the length is 11 feet, find the width of the rectangle.
Answer by Hanifah(3)

Perimeter = 2length +2width
32 = 2.11+2width
32 = 22 +2width
32-22 = 2width
11 = 2 width
11:2 = width
5.5 = width